If ​P(B)=0.2​,​P(A | ​B)=0.7​,​P(B​)=0.8​, and​ P(A|B​)=0.5​, find​ P(B |​ A).
Answer #1

Answer:

Given P(B)=0.2​,​P(A|​B)=0.7​,​P(B'​)=0.8​, and​ P(A|B'​)=0.5
P(A)= P(A∩S)
= P[A∩(BUB')]
= P[(A∩B) U (A∩B')]
= P(A∩B) + P(A∩B')
= P(A|B)P(B) + P(A|B')P(B')
= (0.7)(0.2) + (0.5)(0.8)
= 0.54

P(B|A) = P(A∩B)/P(A)
= 0.7*0.2/0.54
=0.2592